Qd=1500-25P

And Marginal revenue curve by

MR= 20-Q/25

Calculate the perfectly competitive and monopoly outputs and prices                       


1
Expert's answer
2020-08-02T15:34:15-0400

The perfectly competitive outputs and prices is:


"MR=p"

"p=60-\frac {Q}{25}"


"60-\frac {Q}{25}=20-\frac {Q}{25}"

This enterprise cannot operate in perfect competition under the given conditions.


In the case of a monopoly:



"TR=pQ"


"TR=60Q- \frac {Q^2}{25}"


"TR=\int MR"


"TR=20Q-\frac {Q^2}{25}"


"Q=2000"

In the received conditions, the price will be negative.


It is recommended to clarify the condition of the MR function.